OpenStudy (rebeccapink6398):

WILL MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!! Marlin's marble jar had 36 red marbles and 64 blue marbles. What was the ratio of red marbles to blue marbles in the jar? A. 1 : 2 B. 9 : 16 C. 2 : 3 D. 3 : 4

OpenStudy (rebeccapink6398):

PLZ HELP @ms_instrumental :)

OpenStudy (rock_mit182):

easy one divide :

OpenStudy (rock_mit182):

# of blue marbles/# of red marbles

OpenStudy (rock_mit182):

and there is your answer

OpenStudy (rock_mit182):

im sorry is up side down

OpenStudy (rock_mit182):

# of red marbles/ # of blue marbles

OpenStudy (rock_mit182):

;)

OpenStudy (rebeccapink6398):

...I don't get it...

OpenStudy (mathstudent55):

36 red marbles : 64 blue marbles 36 : 64 You can write it as a fraction. Then reduce the fraction. \(\dfrac{36}{64} \)

OpenStudy (mathstudent55):

Both 36 and 64 are divisible by 4.

OpenStudy (rebeccapink6398):

...I don't get get..

OpenStudy (rock_mit182):

the ratio is divided the small number (numerator) by the big number(denominator)

OpenStudy (rock_mit182):

so what is the answer ?

OpenStudy (rebeccapink6398):

..???

OpenStudy (rebeccapink6398):

...I seriously don't get it..

OpenStudy (rock_mit182):



OpenStudy (rock_mit182):

according to the question are more red one's than blue one´s ?

OpenStudy (rebeccapink6398):

ok

OpenStudy (rock_mit182):

once you put the small number in the numerator , and the big in the denominator

OpenStudy (rock_mit182):

just simplify the fraction, that's all

OpenStudy (anonymous):

what is the answer i need help

OpenStudy (mathstudent55):

\(\dfrac{36}{64} = \dfrac{9 \times 4}{16 \times 4} = \dfrac{9}{16} \times \dfrac{4}{4} = \dfrac{9}{16} \times 1 = \dfrac{9}{16} = 9 : 16\)